Can someone help?
The length of the longest leg of a triangle is twice the length of the shortest leg. The remaining leg is 2 inches longer than the shortest leg. If the perimeter of the triangle is 26 inches, how long is each leg?

Answer by Earlsdon(6294)

Let's call the longest leg A and the shortest leg C, so the middle leg will be B.
Now, from the problem, we have:
A = 2C and:
B = C+2 and:
A + B + C = 26 inches. So, if we substitute 2C for A and C+2 for B, we can solve for C.
2C + C+2 + C = 26 Simplify.
4C + 2 = 26 Subtract 2 from both sides of the equation.
4C = 24 Divide both sides by 4.
C = 6 Now that we have C, we can find A and B.
A = 2C = 2(6) = 12
B = C+2 = 6+2 = 8
Answer:
The longest leg is 12 inches.
The shortest leg is 6 inches.
The remaining leg is 8 inches.